OpenStudy (torialaw192):

The length of a rectangle is 24 units. Can the perimeter x of the rectangle be 60 units when its width y is 11 units?

OpenStudy (torialaw192):

Medal for help and best answer! ^^

OpenStudy (torialaw192):

@supercalifragalisticexspeaalli Something wrong do you need the answers??

OpenStudy (supercalifragalisticexspeaalli):

lol is it multiple choice or

OpenStudy (torialaw192):

Yea

OpenStudy (torialaw192):

No, the rectangle cannot have x = 60 and y = 11 because x = 48 + 2y No, the rectangle cannot have x = 60 and y = 11 because x = 24 + 2y Yes, the rectangle can have x = 60 and y = 11 because x + y is greater than 48 Yes, the rectangle can have x = 60 and y = 11 because x + y is greater than 24

OpenStudy (jadeishere):

Okay, so, the perimeter of a rectangle is the added length of all of the sides l + l + w + w = p 24 + 24 + 11 + 11 = 60(?)

OpenStudy (torialaw192):

Right.

OpenStudy (jadeishere):

So, when it's all added together, does it equal 60?

OpenStudy (torialaw192):

No, it equal 70.

OpenStudy (jadeishere):

Okay, so look at the answer options given, there are two for 'no' which one best explains why the perimeter cannot be 60?

OpenStudy (torialaw192):

I think a

OpenStudy (jadeishere):

No xD Okay, so we just discussed that 24 + 24 + 11 + 11 does not equal 60, right? The choice 'a' implies that all those numbers do, in fact, add to 60. Would 24 + 2(11)=60?

OpenStudy (torialaw192):

No. Maybe.

OpenStudy (torialaw192):

No wait, its b.

OpenStudy (jadeishere):

Exactly! Great job

OpenStudy (torialaw192):

Thank you! ^^ Here's a medal for you.

OpenStudy (jadeishere):

Oh, thanks! xD